Carbon dioxide gas is collected at 3.0°C in an evacuated flask with a measured volume of 5.0L. When all the gas has been collected, the pressure in the flask is measured to be 0.090atm . Calculate the mass and number of moles of carbon dioxide gas that were collected. Be sure your answer has the correct number of significant digits.
Answer #1

vapor pressure of water at 30C = 0.0075atm

atmospheric pressure = 0.09atm

The partial pressure of CO2    = 0.09-0.0075 = 0.0825atm

T   = 3+ 273 = 276K

V   = 5L

P   = 0.0825atm

R   = 0.0821L-atm/mole-K

PV = nRT

n   = PV/RT

    = 0.0825atm*5L/(0.0821L-atm/mole-K*276K)

    = 0.018moles of CO2 >>>>.answer